OpenStudy (anonymous):

Use the function value and constraint to evaluate the given trigonometric function: -Function value: sec theta= -4 -constraint: tan theta < 0 -evaluate: cot theta please show work.

OpenStudy (anonymous):

Okay you would need \( \tan^2 \theta +1 = \sec ^2 \theta \)

OpenStudy (anonymous):

and \( \tan \theta = \frac 1 {\cot \theta} \) can you do the rest ?

OpenStudy (anonymous):

I'm confused about the tan squared and sec squared

OpenStudy (anonymous):

It's an identity, try to prove it using right-angle triangle, it's actually very easy.

OpenStudy (anonymous):

but idk what x or y is to get the legs of the right angle triangle

OpenStudy (anonymous):

sec theta = -4 ( r/x) so can sec be -59/14.8?